# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v2

@@ VII Conclusions and outlook
In this paper, we introduce the X3Z3 Floquet code, the first bias-tailored dynamical code based on two-qubit parity check measurements. We show that, despite having no constant stabilisers, the code possesses a persistent symmetry under pure dephasing (or pure bit-flip) noise which allows for a simplified decoding. This results in a substantially improved threshold and sub-threshold performance under biased noise, when compared to other Floquet codes. We demonstrate the enhanced performance through our simulation results obtained from using a fast matching decoder and two error models: a simplistic code-capacity noise model and a noise model approaching realistic circuit-level errors. Besides the superior performance of the X3Z3 Floquet code, our results also show that there are differences, albeit modest, in the threshold and subthreshold performance between the XYZ2 and P6 honeycomb codes under biased noise.
@@
To explain why the X3Z3 Floquet code does not reach the same high performance as the bias-tailored static codes, we prove that a dynamical code on the honeycomb lattice (obeying certain assumptions common to standard Floquet codes) cannot have a so-called 1D-decodable $Z$-detector graph, the crucial requirement for the high performance of static codes. Despite this limitation, the bias-tailored X3Z3 Floquet code has the advantage over its static counterpart in that it requires only lower-weight measurements. Specifically for devices without native two-qubit parity measurements, we devise bias-preserving parity check measurement circuits for any qubit architecture, which allow for high-performance implementation of the code. Our work therefore demonstrates that the X3Z3 Floquet code is a leading quantum error correction code especially for devices with limited connectivity such as the hexagonal and heavy-hexagonal architectures.
@@
We now give several directions for future work. While in this paper we use multi-Pauli product gates for the parity check measurements which are native to certain architectures (for example, Majorana [56, 10, 57] and photonic [58, 17] platforms) we leave for future work the study of the code performance using the bias-preserving parity measurement circuits proposed in Sec. III.2. Furthermore, it would also be interesting to investigate the performance of the X3Z3 Floquet code using a better decoder (such as a tensor network decoder) and to consider ways to analytically derive the best achievable thresholds for the X3Z3 Floquet code.
Author Contributions
Acknowledgements.
We would like to thank our colleagues at Riverlane, especially Alexandra Moylett, Ophelia Crawford, and György Gehér for helpful discussions. We are also grateful to Ophelia Crawford and Hari Krovi for their comments on the manuscript. CM would like to thank Benjamin Brown for pointing out a useful circuit for performing next-nearest-neighbour entangling gates.
Appendix A Details of honeycomb and CSS Floquet codes
A.1 Honeycomb code
We here provide details of the two honeycomb code variants discussed in the main text. The codes’ logical operators evolve through their measurement cycles. In Fig. 10(a), we show one set of logical operators evolving through two measurement cycles. While the measurement sequence has a period of 3, the logical operators only return to their initial values (up to signs) with a period of 6.
In Fig. 10(b), we show the plaquette and edge operators for both honeycomb code variants. The P6 code [8] has edge operators with Pauli bases defined by the edge colour, while the XYZ2 code [6] has edge operators with bases defined by the edge’s orientation in the T junction, i.e., horizontal for the $Z$ edges, and rotated 90 clockwise and counter-clockwise from the horizontal edges, for the $X$ and $Y$ edges, respectively [see Fig. 10(b)]. The plaquette operators of the P6 code also differ between plaquettes of different colours, while the plaquette operators of the XYZ2 code are all the same. Each plaquette operator is a product of the edge operators along the plaquette boundary. For example, a red plaquette operator is the product of the green and blue edge operators around its boundary. The plaquette operators are measured in two subrounds. For instance, after measuring the $g$ and $b$ checks sequentially, we can learn the value of each red plaquette by multiplying the measurement values of the green and blue checks around its boundary. Unlike the CSS and X3Z3 Floquet codes, the honeycomb code has constant plaquette stabilisers.
Figure 10: The honeycomb code measurement cycle, logical operators, plaquettes and edges. (a) A distance-4 (under code-capacity noise) P6 code [8] is shown throughout its evolution. The type of check measured at each measurement subround is indicated by the arrows and the highlighted/coloured edges in the lattice. One set of anti-commuting logical operators is shown in yellow and brown-coloured paths. Each logical is a product of Pauli operators on the coloured qubits along its cycle, where red, green and blue qubits support $X$, $Y$, and $Z$ Pauli bases, respectively. The other set of logical operators is offset by three measurement subrounds. (b) Top: Red, green and blue edges/plaquettes of the P6 honeycomb code support $X$, $Y$ and $Z$ operators, respectively. Bottom: All plaquettes of the XYZ2 honeycomb code support the same XYZ2 operator, as shown on the left. Edge operators are defined based on their orientation in the T junction (rather than the edge colour), as shown on the right.
A.2 CSS Floquet code
In Fig. 11, we illustrate the measurement cycle of the CSS Floquet code [18, 13] and the resulting evolution of (one set of) its logical operators. Its logical operators are either of $Z$-type or $X$-type for all time, as shown. We show a $Z$-type logical operator defined on a vertical non-trivial cycle of the lattice. However, there exists another $X$-type logical operator (not shown) defined on that same vertical cycle. One can find this operator (and its anti-commuting partner) by noting the symmetry in the code obtained by proceeding forward by three measurement subrounds ($t\mapsto t+3$) and applying a Hadamard gate to all qubits ($X\leftrightarrow Z$).
The plaquette operator members of the ISG at each step of the measurement cycle are indicated in Fig. 11. The value of each plaquette operator is inferred, in one time step, from the product of the measurement values of three edges surrounding the plaquette. At each measurement subround, only one kind [which is of one type ($X$ or $Z$) and one colour] of plaquette is reinitialised and one other kind is re-measured to form a detector. The reinitialised and re-measured plaquettes have the same ($X$ or $Z$) type as the checks performed at that particular subround, but are of different colours from each other and also from the checks [18, 13]. The colours of the reinitiliased and re-measured plaquettes are respectively the same as the colours of the checks performed just before and immediately after the current subround.
To see how detectors are formed in this code, let us follow a single blue $X$ plaquette through the measurement sequence. After red $X$ checks (top left of Fig. 11), the blue $X$ plaquette operator is measured (we infer its outcome by multiplying together the check measurement outcomes around its boundary). This initialises the plaquette. It is only re-measured after the green $X$ check measurements, four subrounds later, at which point a detector is formed. After this, blue $Z$ checks are measured, which anti-commute with the blue $X$ plaquette operator. Hence, its measurement outcome after this subround is random.
Figure 11: Measurement cycle for the CSS Floquet code. $X$ or $Z$ edge operators are measured alternately, with the colour of edge and Pauli type indicated by the labelled arrows. Plaquettes labelled with $X$ or $Z$ labels host only a stabiliser operator of the corresponding type, while other plaquettes host both $X$- and $Z$-type stabilisers. One set of anti-commuting logical operators is shown by yellow and brown paths. Each logical is a product of only one type of Pauli operators, either $X$ or $Z$, represented by the red or blue coloured qubits.
Appendix B Noise models
Figure 12: Logical failure probability $p_{L}$ of the X3Z3 Floquet code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}-10^{7}$ number of shots.
@@ Appendix B Noise models
Besides the code-capacity noise model, in this paper we also use the SDEM3 noise model which assumes noisy multi-Pauli product (MPP) measurement gates. The SDEM3 noise model applies single-qubit noise after every single-qubit gate, reset and measurement, two-qubit noise after each MPP gate, and a random classical flip after every measurement. We assume each of the error channels occurs with a total probability $p$, as in Ref. [9]. To take into account biased noise, we generalise the SDEM3 model for depolarising noise used in Refs. [9, 15]. Here the single-qubit noise channel has the same form as given in Eq. (2) for the code-capacity noise model.
Figure 13: Logical failure probability $p_{L}$ of the CSS Floquet code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}$ number of shots.
On top of the single-qubit errors, this model also has a two-qubit noise channel applied after each of the MPP gates, which is given by
@@
Note that Eq. (B) is defined such that the total probability of all the Pauli errors $\sum_{O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}}p_{O}=p$ for all noise biases $\eta$. Apart from the single- and two-qubit Pauli noise channels, we also apply a classical flip of the measurement results with probability $p$ after each of the single and two-qubit measurements. These measurement flips are uncorrelated with the single and two-qubit noise channels.
Appendix C Details of numerical simulations
Figure 14: Logical failure probability $p_{L}$ of the P6 honeycomb code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}$ number of shots.
@@ Appendix C Details of numerical simulations
For each distance $d_{\mathrm{eff}}=d/2$ (where $d_{\mathrm{eff}}$ and $d$ are the minimum number of fault locations required to produce a logical error for the SDEM3 and code-capacity depolarising noise models, respectively), we construct a $2d_{\mathrm{eff}}\times 3d_{\mathrm{eff}}$ honeycomb lattice with periodic boundary conditions. We run quantum memory experiments for $3d_{\mathrm{eff}}$ QEC rounds or $18d_{\mathrm{eff}}$ measurement subrounds (1 QEC round consists of 6 measurement subrounds). The simulations are performed for four different Floquet codes using different values of physical error rates $p$ and various strengths of noise bias $\eta$. Moreover, we consider two different noise models: code-capacity and SDEM3 noise models as detailed in Appendix B.
@@
We choose a single logical qubit for each Floquet code (they all encode two logical qubits on a lattice with periodic boundary conditions) to test for the logical failure probability. Depending on the codes and parameter regimes, we run different numbers of Monte Carlo shots ranging from $10^{5}-10^{10}$ (larger number of shots for longer code distances and smaller physical error rates) for each of the horizontal and vertical logical observables, which give us the horizontal $p_{H}$ and vertical $p_{V}$ logical failure probabilities. We then report the combined logical error probability:
@@
which is an estimate of the probability that either a vertical or a horizontal logical error occurs. Equation (6) assumes that the horizontal and vertical logical errors occur independently. For the infinitely biased code-capacity noise model, one of the logical errors which is of pure $Z$-type or pure $X$-type has zero logical error probability, because the pure biased noise cannot form the other logical operator which is of mixed-type and hence cannot flip the pure-type logical operator. As a result, for the infinitely biased code-capacity noise, the maximum combined logical error probability is $p_{L}=0.5$ [Eq. (6)]. This is in contrast to other noise biases and models where the maximum combined logical error probability is $p_{L}=0.75$ since $p_{V},p_{H}\leq 0.5$.
C.1 Thresholds
Figure 15: Logical failure probability $p_{L}$ of the XYZ2 honeycomb code as a function of $x=(p-p_{\mathrm{th}})d^{1/\nu}$ or $x=(p-p_{\mathrm{th}})d_{\mathrm{eff}}^{1/\nu}$ for two different noise models: Code capacity (upper panels) and SDEM3 (lower panels), with various bias strengths: (a,e) $\eta=1$, (b,f) $\eta=9$, (c,g) $\eta=99$, and (d,h) $\eta=999$. We fit the results to the function $A+Bx+Cx^{2}$ (blue line). Insets: Logical failure probability $p_{L}$ vs physical error rate $p$; thresholds $p_{\mathrm{th}}$ are shown by the magenta vertical lines. Results for each distance $d$ and physical error rate $p$ are averaged over $10^{6}$ number of shots.

### v3

@@ III Discussion
In this paper, we introduce the X3Z3 Floquet code, the first bias-tailored dynamical code based on two-qubit parity check measurements. We show that, despite having no constant stabilisers, the code possesses a persistent stabiliser-product symmetry under pure dephasing (or pure bit-flip) noise which allows for a simplified decoding. This results in a substantially improved threshold and sub-threshold performance under biased noise, when compared to other Floquet codes. We demonstrate the enhanced performance through our simulation results obtained from using a fast matching decoder and two error models: a simplistic code-capacity noise model and a noise model approaching realistic circuit-level errors. Besides the superior performance of the X3Z3 Floquet code, our results also show that there are differences, albeit modest ones, in the threshold and subthreshold performance between the XYZ2 and P6 honeycomb codes under biased noise.
@@
To explain why the X3Z3 Floquet code does not reach the same high performance as the bias-tailored static codes, we prove that a dynamical code on the honeycomb lattice (obeying certain assumptions common to standard Floquet codes) cannot have a 1D decoding graph, the crucial requirement for the high performance of static codes. Despite this limitation, the bias-tailored X3Z3 Floquet code has the advantage over its static counterpart in that it requires only lower-weight measurements. Specifically for devices without native two-qubit parity measurements, we devise phase-bias preserving parity check measurement circuits for any qubit architecture, which allow for high-performance implementation of the code. Our work therefore demonstrates that the X3Z3 Floquet code is a leading quantum error correction code especially for devices with limited connectivity such as the hexagonal and heavy-hexagonal architectures.
@@
We now give several directions for future work. Since the decoding of any general Floquet codes with two-qubit parity measurements can be reduced, at most, to a 2D decoding problem, it would be interesting to investigate the performance improvements of the X3Z3 Floquet code using a more accurate decoder (such as a tensor network decoder) and to consider ways to analytically derive the best achievable thresholds for the X3Z3 Floquet code. Future studies may also investigate fault-tolerant logic in the X3Z3 code, via lattice surgery [7], aperiodic measurement sequences [21], twist braiding [71], or transversal gates [72]. Finally, it will be of interest to study the performance of memory experiments and fault-tolerant logical gate implementations in hardware where parity checks are implemented using bias-preserving circuits such as the circuits presented in this paper and compare them to the performance in hardware with native direct multi-Pauli product measurements.
IV Methods
IV.1 Noise models
@@
We note that, since for Floquet codes one QEC round consists of several measurement subrounds, the single-qubit noise channel in the code-capacity model is applied several times in one QEC round instead of just once as is the case with static codes. The code-capacity noise model is often used as a preferred initial noise model to study before going to a more involved model as its simplicity often offers insight into understanding the code performance.
Noise model Generalised SDEM3
$M_{PP}(p,\eta)$
Noisy gate set Init($p$,$\eta$)
$M(p)$
Measurement ancilla None
Measurement subrounds per one QEC round 6
Table 2: Description of the generalised SDEM3 noise model. The definition of the noisy gates is given in Table 3.
Noisy gate set Description
$M_{PP}(p,\eta)$ Measurement of Pauli Product $PP$ on a pair of qubits: independently register incorrect measurement result with probability $p$, apply a two-qubit $Z$-biased noise channel with total error probability $p$ and noise bias $\eta$ after measurement, i.e., apply: the trivial identity operator ($II$) with a probability $1-p$, $ZI$, $IZ$, and $ZZ$ operators, each with probability $\zeta p/3$, the remaining Pauli operators, each with probability $(1-\zeta)p/12$, where $\zeta=\frac{3}{5}\left(\frac{\eta}{1+\eta}\right)^{2}+\frac{2}{5}\left(\frac{\eta}{1+\eta}\right)$.
Init($p$,$\eta$) Qubit initialization in some Pauli basis, followed by applying a single-qubit $Z$-biased noise channel with total error probability $p$ and noise bias $\eta$ after the reset, i.e., apply: the trivial identity operator ($I$) with a probability $1-p$, Pauli $Z$ operator with probability $p\eta/(\eta+1)$, Pauli $X$ and $Y$ operators, each with probability $p/[2(\eta+1)]$.
$M(p)$ Measure the qubit in some Pauli basis and register the incorrect measurement result with probability $p$.
Table 3: Description of noisy gates in Table 2.
On the other hand, the SDEM3 model is a more elaborate error model involving single-qubit noise channels after every single-qubit gate, measurement and reset, a two-qubit noise channel after every multi-Pauli product (MPP) parity measurement gate, and a classical flip after each measurement. As in Ref. [9], we assume that each of these error channels occurs with a total probability $p$. The SDEM3 noise model is therefore close to standard circuit-level noise and would be a more accurate description of a realistic noise channel, particularly in hardware with native two-qubit measurements, for example, Majorana [62, 10, 63] and photonic [64, 17] architectures.
To take into account biased noise, we generalise the SDEM3 model for depolarising noise used in Refs. [9, 15]. Here the single-qubit noise channel has the same form as given in Eq. (5) for the code-capacity noise model. On top of the single-qubit errors, this model also has a two-qubit noise channel applied after each of the MPP gates, which is given by
@@
Note that Eq. (IV.1) is defined such that the total probability of all the Pauli errors $\sum_{O\in\{I,X,Y,Z\}^{\otimes 2}\backslash\{I\otimes I\}}p_{O}=p$ for all noise biases $\eta$. Apart from the single- and two-qubit Pauli noise channels, we also apply a classical flip of the measurement results with probability $p$ after each of the single and two-qubit measurements. These measurement flips are uncorrelated with the single and two-qubit noise channels. We summarise the description of the SDEM3 noise model and its noisy gate set in Tables 2 and 3, respectively.
IV.2 Stabiliser circuits
We simulate the circuits and generate the error syndromes using Stim [73]; example Stim circuits are provided at Ref. [69]. We construct Stim circuits from various quantum operations (resets, measurement gates, etc.) with noise channels associated with each quantum operation (i.e., with error probabilities set by the noise models introduced above), detectors and logical observable updates. From these circuits, Stim can generate the detector error models which list the error mechanisms, the associated syndromes, and the logical observables flipped by the errors. A detector error model is fed to the decoder which then predicts the most likely error based on a given syndrome. We refer the reader to Ref. [73] for more details on Stim.
IV.3 MWPM decoder
To decode the error syndromes, we apply a minimum-weight perfect matching (MWPM) decoder [74], which is implemented using PyMatching [60]. The syndrome decoding is mapped by the MWPM decoder onto a graph problem, which is subsequently solved by utilizing Edmonds’ algorithm [75, 76] for finding a perfect matching which has minimal weight, i.e., finding a minimum-total-weight set of edges, for which every vertex is incident to exactly one edge [77].
IV.4 Details of numerical simulations
@@
Our simulations are performed for four different Floquet codes using different values of physical error rates $p$ and various strengths of noise bias $\eta$. Moreover, we consider two different noise models: code-capacity and SDEM3 noise models. For each code, the simulations are run with effective distances $d_{\mathrm{eff}}\equiv d/2=4,6,8,10,12$, where the effective distance $d_{\mathrm{eff}}$ is defined as the minimum number of faults under SDEM3 depolarising noise that produce a logical error. This is half of the distance $d$ of the code-capacity noise model. The calculated effective distance differs between noise models since in the SDEM3 noise model, two-qubit errors occurring after an MPP parity gate count as a single fault, while under the code-capacity noise model they would count as two faults.
We choose lattices with periodic boundary conditions. For most simulations, we choose lattice sizes $L\times 3L/2$, where $L=d=2d_{\mathrm{eff}}$ is the lattice length which varies with respect to the code distance. We choose lattice sizes $L\times 3L/2$ such that the horizontal and vertical code distances are the same, as can be inferred from the minimum weight of the horizontal and vertical logical operators (see e.g., Fig. 1(c) for lattice length $L=4$). We simulate the memory experiments for $3d_{\mathrm{eff}}$ QEC rounds or $18d_{\mathrm{eff}}$ measurement subrounds (1 QEC round consists of 6 measurement subrounds). For Supplementary Sec. V of the Supplementary Information [61] where we study elongated codes, we use lattice sizes of $d_{H}\times 3d_{V}/2$ where $d_{V}\geq d_{H}$. Here, $d_{H}$ and $d_{V}$ are the minimum weights of the logical operators in the horizontal and vertical directions, respectively. We simulate the horizontal and vertical logical memory experiments for $23d_{H}/$ 2 and $23d_{V}/$ 2 rounds, respectively.
@@
We select a single logical qubit for each Floquet code (they all encode two logical qubits on a lattice with periodic boundary conditions) to test for the logical failure probability. Depending on the codes and parameter regimes, we run different numbers of Monte Carlo shots ranging from $10^{5}-10^{10}$ (larger number of shots for longer code distances and smaller physical error rates) for each of the horizontal and vertical logical observables, which give us the horizontal $p_{H}$ and vertical $p_{V}$ logical failure probabilities. We then report the combined logical error probability:
@@
which is an estimate of the probability that either a vertical or a horizontal logical error occurs. Equation (9) assumes that the horizontal and vertical logical errors occur independently. For the infinitely phase-biased code-capacity noise model, one of the logical errors which is of pure $Z$-type or pure $X$-type has zero logical error probability, because the pure biased noise cannot form the other logical operator which is of mixed-type and hence cannot flip the pure-type logical operator. As a result, for the infinitely phase-biased code-capacity noise, the maximum combined logical error probability is $p_{L}=0.5$ [Eq. (9)]. This is in contrast to other noise biases and models where the maximum combined logical error probability is $p_{L}=0.75$ since $p_{V},p_{H}\leq 0.5$.
IV.5 Determining the thresholds